### v4

@@
footnotetext: The first author Yang Yue (乐洋) and the sixth author Yang Yue (乐阳) are two distinct individuals sharing the same English name but different Chinese names. Correspond to: {le-y22, zq-chen23}@mails.tsinghua.edu.cn, gaohuang@tsinghua.edu.cn.
Reinforcement Learning with Verifiable Rewards (RLVR) has recently demonstrated notable success in enhancing the reasoning performance of large language models (LLMs), particularly in mathematics and programming tasks. It is widely believed that, similar to how traditional RL helps agents to explore and learn new strategies, RLVR enables LLMs to continuously self-improve, thus acquiring novel reasoning abilities that exceed the capacity of the corresponding base models. In this study, we take a critical look at the current state of RLVR by systematically probing the reasoning capability boundaries of RLVR-trained LLMs across various model families, RL algorithms, and math/coding/visual reasoning benchmarks, using pass@k at large k values as the evaluation metric. While RLVR improves sampling efficiency towards correct paths, we surprisingly find that current training does not elicit fundamentally new reasoning patterns. We observe that while RLVR-trained models outperform their base models at smaller values of $k$ (e.g., kk=1), base models achieve higher pass@$k$ score when $k$ is large. Moreover, we observe that the reasoning capability boundary of LLMs often narrows as RLVR training progresses. Further coverage and perplexity analysis shows that the reasoning paths generated by RLVR models are already included in the base models’ sampling distribution, suggesting that their reasoning abilities originate from and are bounded by the base model. From this perspective, treating the base model as an upper bound, our quantitative analysis shows that six popular RLVR algorithms perform similarly and remain far from optimal in fully leveraging the potential of the base model. In contrast, we find that distillation can introduce new reasoning patterns from the teacher and genuinely expand the model’s reasoning capabilities. Taken together, our findings suggest that current RLVR methods have not fully realized the potential of RL to elicit genuinely novel reasoning abilities in LLMs. This underscores the need for improved RL paradigms–such as continual scaling and multi-turn agent-environment interaction–to unlock this potential.    Project Page: https://limit-of-RLVR.github.io
@@ 1 Introduction
The development of reasoning-centric large language models (LLMs), such as OpenAI-o1 Jaech et al. (2024), DeepSeek-R1 Guo et al. (2025), and Kimi-1.5 Team et al. (2025), has significantly advanced the frontier of LLM capabilities, particularly in solving complex logical tasks involving mathematics and programming. In contrast to traditional instruction-tuned approaches that rely on human-curated annotations Achiam et al. (2023); Grattafiori et al. (2024), the key driver behind this leap forward is large-scale Reinforcement Learning with Verifiable Rewards (RLVR) Lambert et al. (2024); Guo et al. (2025). RLVR starts with a pretrained base model or one fine-tuned on long chains of thought (CoT) data, optimizing it via reinforcement learning based on simple, automatically computable rewards. These rewards are determined by whether the model’s output matches a ground-truth solution in mathematics or passes unit tests in code, thus enabling scalability without human labeling. This framework has gained significant attention due to its simplicity and practical effectiveness. In traditional RL settings such as game playing (e.g., Atari, Go), agents often autonomously discover new strategies and surpass even human-level performance through self-improvement Mnih et al. (2015); Silver et al. (2017). Inspired by this success, it is widely believed that RLVR similarly enables LLMs to autonomously develop novel reasoning patterns, including enumeration, self-reflection, and iterative refinement, surpassing the capabilities of their base models Guo et al. (2025). Consequently, RLVR has been considered a promising path toward continuously self-evolving LLMs, potentially bringing us closer to more powerful intelligence Guo et al. (2025).
@@
To rigorously answer this question, we must first assess the reasoning capability boundaries of both base and RLVR-trained models. Traditional evaluation metrics rely on average score from greedy decoding or nucleus sampling Holtzman et al. (2020), which reflects average-case behavior. However, these metrics risk underestimating the true potential of a model, especially if it fails on difficult problems after limited attempts, despite being capable of solving them with more sampling. To overcome this limitation, we adopt the pass@k metric Brown et al. (2024), where a problem is considered solved if any of the $k$ sampled outputs is correct. By allowing multiple attempts, pass@k reveals whether a model has the potential to solve a problem. The average pass@k across a dataset thus reflects the proportion of problems a model can potentially solve within $k$ trials, offering a more robust view of its reasoning boundary. This provides a rigorous test on whether the RLVR training yields fundamentally transcending capacity, enabling the model to solve problems that the base model cannot.
@@
$\bullet$ Current RLVR models exhibit narrower reasoning coverage than their base models. In pass@$k$ curves, although RLVR models outperform their base models at small $k$, it is surprising that base models consistently surpass RLVR models across all benchmarks and LLM families as $k$ increases. This suggests that current RLVR training does not expand, and even reduce the scope of reasoning over solvable problems. Manual inspection of model responses shows that, for most problems, the base model can produce at least one correct CoT, implying that it can already generate correct reasoning paths for problems that were previously considered only solvable for RLVR models.
@@
In conclusion, our findings show that current RLVR methods, while improving sampling efficiency, do not elicit novel reasoning beyond the base model’s capabilities. This highlights a gap between existing RLVR methods and the goals of reinforcement learning, underscoring the need for improved RL paradigms such as continual scaling, better exploration, and multi-turn agent interaction.
@@ 2.1 Reinforcement Learning with Verifiable Rewards
RLVR Algorithms. Proximal Policy Optimization (PPO) Schulman et al. (2017) proposed using the following clipped surrogate to maximize the objective:
@@
where $r_{t}(\theta)=\pi_{\theta}(y_{t}|x,\mathbf{y}_{<t})/\pi_{\theta_{\text{old}}}(y_{t}|x,\mathbf{y}_{<t})$, and $A_{t}$ is the advantage estimated by a value network $V_{\phi}$. KL divergence term is optionally applied, to constrain the model from deviating too far from the original policy. More algorithms are introduced in Section˜D.5.
Policy Gradient. PPO and its variants belong to the policy gradient class of RL  Williams (1992); Sutton et al. (1998). These methods learn exclusively from on-policy samples, i.e., samples generated by the current LLM. In the context of verifiable rewards, the training objective generally maximizes the log-likelihood of samples with correct answers and minimizes the likelihood of those with incorrect answers.
Zero RL Training applies RL directly to the base model without any supervised fine-tuning (SFT) Guo et al. (2025). To clearly study the effect of RLVR, we follow this zero-RL setting for all math tasks using pretrained models as start model. However, for coding and visual reasoning tasks, open-source work typically uses instruction-tuned models as starting points, primarily due to the training instability and limited effectiveness of using a pure zero-RL setting. Following this convention, we compare the finetuned model with its RLVR-trained counterpart to focus solely on the effect of RLVR.
@@ 2.2 Metrics for LLM Reasoning Capacity Boundary
Pass@$k$ Metrics. Accurately measuring the reasoning ability boundary of base and RL models is challenging, as methods like greedy decoding or the average of nucleus samplings Holtzman et al. (2020) only reflect average-case performance. To accurately measure the reasoning ability boundary, we extend the commonly used pass@k metric from code generation Chen et al. (2021) to all tasks with verifiable rewards. Given a problem, we sample $k$ outputs from the model. The pass@$k$ value for this question is 1 if at least one of the $k$ samples passes verification; otherwise, it is 0. The average pass@$k$ value over the dataset reflects the proportion of problems in the dataset that the model can solve within $k$ trials, providing a rigorous evaluation of the reasoning capacity coverage of LLMs. We adopt an unbiased, low-variance estimator for computing to calculate pass@$k$, as detailed in Section˜A.2.
Comparison with Best-of-$N$ and Majority Voting. Best-of-$N$ Cobbe et al. (2021) and majority voting are practical methods for selecting correct answers, but they may overlook a model’s full reasoning potential. In contrast, we use pass@$k$ not to assess practical utility but to investigate the boundaries of reasoning capacity. If a model produces a correct solution in any of the $k$ samples, we treat the problem as within its potential scope. Thus, if RL enhances reasoning, the RL-trained model should succeed in more such problems than the base model. Methods like Best-of-$N$ or majority voting may miss these successes if the correct answer is not selected by the verifier or voting.
Random Guessing Issue. For coding tasks, where a compiler and predefined unit test cases are used as verifiers, the pass@$k$ value can accurately reflect whether the model can solve the problem. In mathematics, the issue of “guessing” can become pronounced as $k$ increases, where a model may generate an incorrect CoT but still accidentally arrive at the correct answer. To address this, we manually check the correctness of CoT for a subset of model outputs as detailed in Section˜3.1. By combining reuslts on math with manually checking and coding, we rigorously evaluate the scope of LLM’s reasoning capacity. Another caveat is that, with an astronomically large $k$, even uniform sampling over the token dictionary would stumble upon the correct reasoning path–though this is infeasible within today’s time and compute budgets. Crucially, we find that the base model already produces correct outputs at realistic values ($k=128$ or $1024$), well within practical resource limits.
@@ 3 RLVR’s Effect on Reasoning Capacity Boundary
Evaluation Protocol. For sampling procedures for both base and RLVR models, we use a temperature of 0.6 and a top-$p$ value of 0.95, allowing a maximum generation of 16,384 tokens. We also show the effect of different temperature settings in Figure˜17. For evaluation of the base model, a common practice is to include few-shot examples in the prompt to guide the output Grattafiori et al. (2024); Yang et al. (2024); Liu et al. (2024). However, to ensure a fair and unbiased comparison, we deliberately avoid using few-shot prompts for base models, eliminating any potential confounding effects on reasoning that might be introduced by in-context examples. For evaluating both the base and RLVR models, we use the same zero-shot prompt as in RLVR training, or the default prompt provided by the benchmark, ensuring a consistent setup across both models. Interestingly, although base models often produce unformatted or non-sensical responses without few-shot guidance, we observe that with sufficient sampling, they are still capable of generating correctly formatted outputs and successfully solving complex problems. Prompt templates for training and evaluation are provided in Appendix˜E.
@@ 3.1 RLVR for Mathematical Reasoning
Models and Benchmarks. In math problems, models are required to generate a reasoning process (i.e., CoT) along with the final answer. To ensure the robustness of conclusions, we experiment with multiple LLM families, primarily Qwen2.5 (7B/14B/32B base variants) Yang et al. (2024) and additionally LLaMA-3.1-8B Grattafiori et al. (2024). We adopt RLVR models released by SimpleRLZoo Zeng et al. (2025), which train zero-RL models using GRPO on GSM8K and the MATH training set, with correctness reward only, excluding any format-based reward. We compare the pass@$k$ curves of base and zero-RL models on benchmarks of varying difficulty: GSM8K Cobbe et al. (2021), MATH500 Hendrycks et al. (2021), Minerva Lewkowycz et al. (2022), Olympiad He et al. (2024), AIME24, and AMC23. Additionally, we include the RLVR model Oat-Zero-7B and DAPO-32B Liu et al. (2025b); Yu et al. (2025). These two models are characterized by strong performance on the challenging AIME24 benchmark.
@@
Validity of Chain-of-Thought. For mathematical problems, the common evaluation is based solely on the correctness of the final answer, with the risk of “hacking”. To accurately reflect the reasoning ability boundary using pass@$k$, it is important to assess how many solved problems result from sampling genuinely correct CoTs, rather than from lucky guesses. Following Brown et al. (2024), we manually inspect all CoTs that led to correct answers to the most challenging solvable problems in the GSM8k dataset – those with an average accuracy below 5% but above 0%. The base model answered 25 such questions, with 24 containing at least one correct CoT. Similarly, the RL-trained model answered 25 questions, 23 of which included at least one correct CoT. We also manually check the CoTs for problems in the challenging AIME24 benchmark with an average accuracy below 5%. Details can be found in Section˜D.2. The base model answered 7 such questions, with 5 out of 6 containing at least one correct CoT (excluding one ambiguous case of correctness due to skipped reasoning steps). Similarly, the RL-trained model answered 6 questions, 4 of which included at least one correct CoT. These results suggest that the base model can sample valid reasoning paths to solve the problems.
@@ 3.2 RLVR for Code Generation
Models and Benchmarks. We adopt the open-sourced RLVR-trained model, CodeR1-Zero-Qwen2.5-7B Liu & Zhang (2025), which trains zero-RL models on 12K LeetCode and TACO samples over 832 steps, based on Qwen2.5-7B-Instruct-1M Yang et al. (2025b). For evaluation, models are assessed on LiveCodeBench v5, comprising 279 problems that span from August 2024 to January 2025 Jain et al. (2025), as well as HumanEval+ and MBPP+ Liu et al. (2023). We also evaluate the most powerful open-source RLVR-trained coding LLM, DeepCoder-14B Luo et al. (2025), built on DeepSeek-R1-Distill-Qwen-14B. Here both models take 32k response length. Due to their high computational cost, we evaluate them only on LiveCodeBench as a representative benchmark.

### v5

@@
footnotetext: The first author Yang Yue (乐洋) and the sixth author Yang Yue (乐阳) share the same English name but different Chinese names. Correspond to: {le-y22, zq-chen23}@mails.tsinghua.edu.cn, gaohuang@tsinghua.edu.cn.
Reinforcement Learning with Verifiable Rewards (RLVR) has recently demonstrated notable success in enhancing the reasoning performance of large language models (LLMs), particularly in mathematics and programming tasks. It is widely believed that, similar to how traditional RL helps agents to explore and learn new strategies, RLVR enables LLMs to continuously self-improve, thus acquiring novel reasoning abilities that exceed the capacity of the corresponding base models. In this study, we take a critical look at the current state of RLVR by systematically probing the reasoning capability boundaries of RLVR-trained LLMs across various model families, RL algorithms, and math/coding/visual reasoning benchmarks, using pass@k at large k values as the evaluation metric. While RLVR improves sampling efficiency towards correct paths, we surprisingly find that current training rarely elicit fundamentally new reasoning patterns. We observe that while RLVR-trained models outperform their base models at smaller values of $k$ (e.g., kk=1), base models achieve higher pass@$k$ score when $k$ is large. Moreover, we observe that the reasoning capability boundary of LLMs often narrows as RLVR training progresses. Further coverage and perplexity analysis shows that the reasoning paths generated by RLVR models are already included in the base models’ sampling distribution, suggesting that their reasoning abilities originate from and are bounded by the base model. From this perspective, treating the base model as an upper bound, our quantitative analysis shows that six popular RLVR algorithms perform similarly and remain far from optimal in fully leveraging the potential of the base model. In contrast, we find that distillation can introduce new reasoning patterns from the teacher and genuinely expand the model’s reasoning capabilities. Taken together, our findings suggest that current RLVR methods have not fully realized the potential of RL to elicit genuinely novel reasoning abilities in LLMs. This underscores the need for improved RL paradigms, such as effective exploration mechanism, more deliberate and large-scale data curation, fine-grained process signal, and multi-turn agent interaction, to unlock this potential.    Project Page: https://limit-of-RLVR.github.io
@@ 1 Introduction
The development of reasoning-centric large language models (LLMs), such as OpenAI-o1 jaech2024openaio1, DeepSeek-R1 guo2025deepseek-r1, and Kimi-1.5 team2025kimi, has significantly advanced the frontier of LLM capabilities, particularly in solving complex logical tasks involving mathematics and programming. In contrast to traditional instruction-tuned approaches that rely on human-curated annotations achiam2023gpt4; grattafiori2024llama3, the key driver behind this leap forward is large-scale Reinforcement Learning with Verifiable Rewards (RLVR) lambert2024tulu3; guo2025deepseek-r1. RLVR starts with a pretrained base model or one fine-tuned on long chains of thought (CoT) data, optimizing it via reinforcement learning based on simple, automatically computable rewards. These rewards are determined by whether the model’s output matches a ground-truth solution in mathematics or passes unit tests in code, thus enabling scalability without human labeling. This framework has gained significant attention due to its simplicity and practical effectiveness. In traditional RL settings such as game playing (e.g., Atari, Go), agents often autonomously discover new strategies and surpass even human-level performance through self-improvement mnih2015dqn; silver2017alphagozero. Inspired by this success, it is widely believed that RLVR similarly enables LLMs to autonomously develop novel reasoning patterns, including enumeration, self-reflection, and iterative refinement, surpassing the capabilities of their base models guo2025deepseek-r1. Consequently, RLVR has been considered a promising path toward continuously self-evolving LLMs, potentially bringing us closer to more powerful intelligence guo2025deepseek-r1.
@@
To rigorously answer this question, we must first assess the reasoning capability boundaries of both base and RLVR-trained models. Traditional evaluation metrics rely on average score from greedy decoding or nucleus sampling holtzman2019topp, which reflects average-case behavior. However, these metrics risk underestimating the true potential of a model, especially if it fails on difficult problems after limited attempts, despite being capable of solving them with more sampling. To overcome this limitation, we adopt the pass@k metric brown2024monkey, where a problem is considered solved if any of the $k$ sampled outputs is correct. By allowing multiple attempts, pass@k reveals whether a model has the potential to solve a problem. The average pass@k across a dataset thus reflects the proportion of problems a model can potentially solve within $k$ trials, offering a more robust view of its reasoning boundary. This provides a rigorous test on whether the RLVR training yields fundamentally transcending capacity, enabling the model to solve problems that the base model cannot.
@@
$\bullet$ Current RLVR models often exhibit narrower reasoning coverage than their base models. In pass@$k$ curves, although RLVR models outperform their base models at small $k$, it is surprising that base models consistently surpass RLVR models across all benchmarks and LLM families as $k$ increases. This suggests that current RLVR training does not expand, and even reduce the scope of reasoning over solvable problems. Manual inspection of model responses shows that, for most problems, the base model can produce at least one correct CoT, implying that it can already generate correct reasoning paths for problems that were previously considered only solvable for RLVR models.
@@
In conclusion, our findings show that current RLVR methods, while improving sampling efficiency, rarely elicit novel reasoning beyond the base model. This highlights a gap between existing RLVR methods and the goals of reinforcement learning, underscoring the need for improved RL paradigms such as better exploration, continual data scaling, fine-grained process signal, and multi-turn agent interaction.
@@ 2.1 Reinforcement Learning with Verifiable Rewards
RLVR Algorithms. Proximal Policy Optimization (PPO) schulman2017ppo proposed using the following clipped surrogate to maximize the objective:
@@
where $r_{t}(\theta)=\pi_{\theta}(y_{t}|x,\mathbf{y}_{<t})/\pi_{\theta_{\text{old}}}(y_{t}|x,\mathbf{y}_{<t})$, and $A_{t}$ is the advantage estimated by a value network $V_{\phi}$. KL divergence term is optionally applied, to constrain the model from deviating too far from the original policy. More algorithms are introduced in Section˜C.5.
Policy Gradient. PPO and its variants belong to the policy gradient class of RL  williams1992reinforce; sutton1998reinforcement. These methods learn exclusively from on-policy samples, i.e., samples generated by the current LLM. In the context of verifiable rewards, the training objective generally maximizes the log-likelihood of samples with correct answers and minimizes the likelihood of those with incorrect answers.
Zero RL Training applies RL directly to the base model without any supervised fine-tuning (SFT) guo2025deepseek-r1. To clearly study the effect of RLVR, we follow this zero-RL setting for all math tasks using pretrained models as start model. However, for coding and visual reasoning tasks, open-source work typically uses instruction-tuned models as starting points, primarily due to the training instability and limited effectiveness of using a pure zero-RL setting. Following this convention, we compare the finetuned model with its RLVR-trained counterpart to focus solely on the effect of RLVR.
@@ 2.2 Metrics for LLM Reasoning Capacity Boundary
Pass@$k$ Metrics. Accurately measuring the reasoning ability boundary of base and RL models is challenging, as methods like greedy decoding or the average of nucleus samplings holtzman2019topp only reflect average-case performance. To accurately measure the reasoning ability boundary, we extend the commonly used pass@k metric from code generation chen2021evaluating to all tasks with verifiable rewards. Given a problem, we sample $k$ outputs from the model. The pass@$k$ value for this question is 1 if at least one of the $k$ samples passes verification; otherwise, it is 0. The average pass@$k$ value over the dataset reflects the proportion of problems in the dataset that the model can solve within $k$ trials, providing a rigorous evaluation of the reasoning capacity coverage of LLMs. We adopt an unbiased, low-variance estimator for computing to calculate pass@$k$, as detailed in Section˜A.2.
Comparison with Best-of-$N$ and Majority Voting. Best-of-$N$ cobbe2021gsm8k and majority voting are practical methods for selecting correct answers, but they may overlook a model’s full reasoning potential. In contrast, we use pass@$k$ not to assess practical utility but to investigate the boundaries of reasoning capacity. If a model produces a correct solution in any of the $k$ samples, we treat the problem as within its potential scope. Thus, if RL enhances reasoning, the RL-trained model should succeed in more such problems than the base model. Methods like Best-of-$N$ or majority voting may miss these successes if the correct answer is not selected by the verifier or voting.
Random Guessing Issue. For coding tasks, where a compiler and predefined unit test cases are used as verifiers, the pass@$k$ value can accurately reflect whether the model can solve the problem. In mathematics, the issue of “guessing” can become pronounced as $k$ increases, where a model may generate an incorrect CoT but still accidentally arrive at the correct answer. To address this, we manually check the correctness of CoT for a subset of model outputs as detailed in Section˜3.1. By combining results on math with manually checking and coding, we rigorously evaluate the scope of LLM’s reasoning capacity. Another caveat is that, with an astronomically large $k$, even uniform sampling over the token dictionary would stumble upon the correct reasoning path–though this is infeasible within today’s time and compute budgets. Crucially, we find that the base model already produces correct outputs at realistic values ($k=128$ or $1024$), well within practical resource limits.
@@ 3 RLVR’s Effect on Reasoning Capacity Boundary
Evaluation Protocol. For sampling procedures for both base and RLVR models, we use a temperature of 0.6 and a top-$p$ value of 0.95, allowing a maximum generation of 16,384 tokens. We also show the effect of different temperature settings in Figure˜17. For evaluation of the base model, a common practice is to include few-shot examples in the prompt to guide the output grattafiori2024llama3; yang2024qwen25; liu2024deepseekv3. However, to ensure a fair and unbiased comparison, we deliberately avoid using few-shot prompts for base models, eliminating any potential confounding effects on reasoning that might be introduced by in-context examples. For evaluating both the base and RLVR models, we use the same zero-shot prompt as in RLVR training, or the default prompt provided by the benchmark, ensuring a consistent setup across both models. Interestingly, although base models often produce unformatted or non-sensical responses without few-shot guidance, we observe that with sufficient sampling, they are still capable of generating correctly formatted outputs and successfully solving complex problems. Prompt templates for training and evaluation are provided in Appendix˜D.
@@ 3.1 RLVR for Mathematical Reasoning
Models and Benchmarks. In math problems, models are required to generate a reasoning process (i.e., CoT) along with the final answer. To ensure the robustness of conclusions, we experiment with multiple LLM families, primarily Qwen2.5 (7B/14B/32B base variants) yang2024qwen25 and additionally LLaMA-3.1-8B grattafiori2024llama3. We adopt RLVR models released by SimpleRLZoo zeng2025simplerl, which train zero-RL models using GRPO on GSM8K and the MATH training set, with correctness reward only, excluding any format-based reward. We compare the pass@$k$ curves of base and zero-RL models on benchmarks of varying difficulty: GSM8K cobbe2021gsm8k, MATH500 hendrycks2021math500, Minerva lewkowycz2022minerva, Olympiad he2024olympiadbench, AIME24, and AMC23. Additionally, we include the RLVR model Oat-Zero-7B and DAPO-32B liu2025oat; yu2025dapo. These two models are characterized by strong performance on the challenging AIME24 benchmark.
@@
Validity of Chain-of-Thought. For mathematical problems, the common evaluation is based solely on the correctness of the final answer, with the risk of “hacking”. To accurately reflect the reasoning ability boundary using pass@$k$, it is important to assess how many solved problems result from sampling genuinely correct CoTs, rather than from lucky guesses. Following brown2024monkey, we manually inspect all CoTs that led to correct answers to the most challenging solvable problems in the GSM8k dataset – those with an average accuracy below 5% but above 0%. The base model answered 25 such questions, with 24 containing at least one correct CoT. Similarly, the RL-trained model answered 25 questions, 23 of which included at least one correct CoT. We also manually check the CoTs for problems in the challenging AIME24 benchmark with an average accuracy below 5%. Details can be found in Section˜C.2. The base model answered 7 such questions, with 5 out of 6 containing at least one correct CoT (excluding one ambiguous case of correctness due to skipped reasoning steps). Similarly, the RL-trained model answered 6 questions, 4 of which included at least one correct CoT. These results suggest that the base model can sample valid reasoning paths to solve the problems.
@@ 3.2 RLVR for Code Generation
Models and Benchmarks. We adopt the open-sourced RLVR-trained model, CodeR1-Zero-Qwen2.5-7B code-r1, which trains zero-RL models on 12K LeetCode and TACO samples over 832 steps, based on Qwen2.5-7B-Instruct-1M yang2025qwen25-1m. For evaluation, models are assessed on LiveCodeBench v5, comprising 279 problems that span from August 2024 to January 2025 jain2024livecodebench, as well as HumanEval+ and MBPP+ evalplus. We also evaluate the most powerful open-source RLVR-trained coding LLM, DeepCoder-14B deepcoder2025, built on DeepSeek-R1-Distill-Qwen-14B. Here both models take 32k response length. Due to their high computational cost, we evaluate them only on LiveCodeBench as a representative benchmark.